OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 150r2 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r cubed square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (anonymous):

to get the number of papers, we have to divide the whole area with area of one piece of paper. area/piece of paper \[\frac{ 150r ^{2} }{ \sqrt{r ^{3}} }=150*\frac{ r ^{2} }{ r ^{\frac{ 3 }{ 2 }} }\] now you just divide r^2 with r^(3/2) and that is the area you are looking for;) if you have any question or troubles with the division, feel free to ask